OpenStudy (anonymous):

whats the area of a quarter-circle with diameter of 4 meters

OpenStudy (anonymous):

90/360 * pi *2^2

OpenStudy (anonymous):

as radius is half of diameter solve it!

OpenStudy (saifoo.khan):

You are looking for this area.

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

so the formula for a full circle is A = pi r² so if we have a quarter all we need is 1/4 of that so we just do A = 1/4 pi r² we have 4m for the diameter and the radius is 1/2 that or 2 so we plug in 2 into r A = 1/4 pi 2^2 A = 1/4 pi 4 now we get 1/4 of 4 so we end up with 1 A = 1pi m^2

OpenStudy (anonymous):

pi is equal to 3.14 so we can just say A = 3.14 m^2